Differenciate with respect to x, when y=1+cos2x/1+sinx^

Grade:12th pass

2 Answers

Arun
5 years ago
Dear Prakash
 
Please complete your question as in the RHS expression, it is not complete.
 
Please check and repost it. I will be happy to help you
Nitesh kumar
5 years ago
Y=(1+cos2x)/(1+sinx)
Y=(1+2cos2x-1)/(1+sinx)
Y=2cos2x/(1+sinx)
Y=2(1-sin2x)/(1+sinx)
Y=2(1+sinx)(1-sinx)/(1+sinx)
Y=2(1+sinx)
Now differentiate wrt x
dy/dx=2d(1+sinx)/dx
dy/dx=2cosx